OpenStudy (anonymous):

What is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x + y = 1 y = 6 - 4x -x + 6 = 1 x + 6 = 1 7x + 6 = 1

OpenStudy (anonymous):

wait, so these: -x + 6 = 1 x + 6 = 1 7x + 6 = 1 are the answer choices?